# V2X Cooperative Perception for Autonomous Driving: Recent Advances and Challenges

### v3

@@ IV-B1 Cooperative Agent Selection
Selecting the right collaborators is crucial because inappropriate candidates can cause disruptions and disturbances that could harm overall performance. In recent years, communication strategies have been created to help choose beneficial partners for cooperation, considering perception improvement and communication quality.
The attention mechanism, recognized for its efficacy in improving model efficiency by focusing on relevant data, offers promise in selecting suitable agents for the ego CAV. In this context, an attention-based communication scheme was introduced in [150] for V2I cooperative object detection. This scheme leverages a general attention mechanism, where each infrastructure calculates a matching score based on local and received information within their cooperative system. Subsequently, they broadcast this score along with CPMs to the collaborative vehicle. Upon receiving data packets from connected infrastructures, the ego vehicle employs an attention mechanism to select a neighbouring infrastructure for cooperation. While this approach conserves transmission bandwidth by choosing only one agent for V2I communication, it does not harness global information optimally.
Multi-agent Cooperation can be modelled as a topology graph where each node represents a connected agent and transmission messages are generated and broadcasted periodically. In [151], a graph-based on-demand data sharing mechanism was proposed to address network constraints. The feature maps from each vehicle are compressed through quantization to reduce communication costs. Ego-vehicles receive feature maps only from selected coop-vehicles located within their perception range, establishing communication channels and continuously receiving the message body. The on-demand sharing mechanism enabled efficient cooperation among connected vehicles by selectively sharing compressed feature maps but ignored the integration of spatial-temporal information.
@@
To minimize energy consumption, the sensor sharing scheduling problem can be framed as a variant of the Multi-Armed Bandit (MAB) problem [175], factoring in perception performance, time-varying wireless channels, and power consumption. In this regard, an online learning-based algorithm known as the Adaptive Volatile Upper Confidence Bound (AVUCB) algorithm was proposed by reference [152]. This algorithm schedules the most advantageous vehicle, offering an enhanced view while minimizing total power consumption and adhering to communication bandwidth and detection accuracy constraints. However, the AVUCB algorithm does not account for the high mobility of cooperative vehicles. Addressing this mobility challenge, the Mobility-Aware Sensor Scheduling (MASS) algorithm [154], an extension of the AVUCB, formulated decentralized CP scheduling as a Restless-MAB (RMAB) problem [176]. In the RMAB context, rewards continually evolve, requiring each vehicle to continuously learn about its surroundings while selecting actions that empirically maximize rewards. When applied to a practical SUMO trace, the MASS algorithm demonstrated the ability to enhance overall perception gain without incurring additional costs associated with frequent meta-information exchanges.
@@
Cooperative agent selection can also be likened to a Multi-agent Path Finding (MAPF) problem, which plans paths for multiple agents [177]. Inspired by MAPF, a selective communication algorithm [155] was proposed to determine the subset of cooperative vehicles based on information gains from connected vehicles. Each vehicle predicts the set of objects in its field of view and transmits the centre locations of all detected objects to the ego-vehicle. Upon receiving this location information from all vehicles, the ego-vehicle estimates information gains by comparing its detected objects’ positions to received 2D centre positions from other connected cars. The vehicles associated with the highest information gain are selected to form the communication scope. In contrast to random selection, this algorithm achieved higher success rates with minimal additional communication costs.
@@
However, all the aforementioned methods primarily focused on spatial correlations, overlooking the significance of temporal dependencies in collaborator selection. To address this gap, IoSI-CP [156] capitalized on the importance of semantic information from both temporal and spatial dimensions to select optimal collaborators. This approach leveraged a Graph Neural Network (GNN) to capture inherent correlations between the ego vehicle and its neighbouring agents. Each agent was represented as a node, encompassing sparse feature maps and latency information. The GNN assigned weight values to each agent with respect to the ego vehicle. Agents with weights of zero or lower were considered ineligible for collaboration. This refined weighting effectively eliminated unsuitable collaborators responsible for noise, leading to improved perception performance. Nevertheless, this approach relied on one-dimensional weights, which may not fully capture the intricacies addressed by high-dimensional representation methods.
@@ IV-B2 Perception Information Selection
Recently, there has been a growing concern regarding the selection of cooperative message content. A number of methods have emerged for selecting or generating CPMs intended for transmission. These methods can be broadly categorized into rule-based, distance-based, and learning-based approaches, each rooted in distinct mechanisms and theoretical principles.
@@
The rule-based approach to CPM generation/selection entails designing rules that dictate which aspects of detected object information should be included in a CPM. For example, the work in [157] proposed a rule that transmits information about an object’s mode (including acceleration and yaw rate) only if the object is turning or accelerating. Similarly, in [158], visual information from detected vehicles, such as license plates, vehicle brands, and colours, was used to determine their communication IDs. These IDs are then used to integrate CPMs with local perception information, ensuring the confidentiality of the V2X architecture. All detected vehicles that have exchanged CAM messages will not be considered in future CPMs. Standardization efforts, led by the European Telecommunications Standards Institute (ETSI), have also introduced CPM generation rules, specifying what information CPMs should include and when to generate new messages. Further, the work in [159] and [160] has proposed improved algorithms for these rules to filter out redundant information about detected objects that nearby vehicles have recently transmitted. While rule-based methods are straightforward to implement, they rely primarily on changes in the position and speed of detected objects.
@@
Distance-based methods prioritize data based on distance metrics, considering the spatial location relative to the ego vehicle’s onboard sensors to measure data point significance. For instance, the Augmented Informative Collaborative Perception (AICP) system [161] employed a fitness sorting algorithm based on Mahalanobis distance to select crucial information at the application level, effectively filtering out less relevant packets with minimal delay. In a separate approach, Dynamic Feature Sharing (DFS) [162] addresses limited communication bandwidth challenges by identifying relevant feature data for sharing through Random-K priorities based on Manhattan distance. DFS incorporates early, intermediate, and late fusion strategies within the CP system to optimize dynamic node engagement.
Rule-based methods primarily select specific attributes of detected objects for dissemination, while distance-based methods emphasize the transmission of feature cells or data points close to the ego vehicle. However, these approaches may only partially meet the ego vehicles’ perception information requirements in real-world scenarios, leading to suboptimal results. In contrast, learning-based approaches leverage the entirety of perception data to intelligently determine information selection for sharing.
@@
Deep Reinforcement Learning (DRL) allows agents to learn policies through interactions with their environment, maximizing cumulative rewards [178, 179, 180]. DRL techniques can be applied to vehicular communications to optimize data selection for transmission. The work [163] introduced a DRL-based Cooperative Perception (DRLCP) scheme, where each CAV intelligently decides what information to transmit based on the context captured by its onboard sensors. DRLCP’s primary objective is to determine whether to transmit or discard data, allocating communication resources to critical packets and reducing redundancy. However, it was evaluated on only two road networks without considering the impact of obstructions on vehicular communications.
Sparse Bird’s Eye View (BEV) features extracted from point clouds can be further compressed to minimize redundancy by extracting valuable semantic information. Feature encoding with Convolutional Neural Networks (CNNs) is a widely adopted learning-based method for reducing CPM size. For instance, FS-COD [164] modified the backbone structure of Complex-YOLO [181] based on BEV image characteristics. Filters at the last layer of the CNN based feature extractor determine the size of shared data between cooperative vehicles. COOPERNAUT [165] designed a Point Encoder using the Point Transformer [182] for each connected vehicle to learn LiDAR information representations. This approach processes and encodes point clouds into key points with compact representations learned by Point Transformer blocks, filtering partially processed data to balance perception accuracy and communication load. CenterCoop [166] encoded the informative cues from the whole BEV context into compact center representations. The communication cost can be significantly reduced by transmitting compact center features as queries instead of the whole feature map.
While CNN-generated feature maps offer potential data volume reduction, their dimensions often exceed transmission feasibility within current V2X technology scopes [167]. To address this concern, a comprehensive strategy was adopted in [167], combining feature encoding, channel selection, and data compression techniques to optimize feature maps. Specifically, an autoencoder architecture, including a channel-wise semantic information encoder and a pointwise convolution decoder, optimized the representation of sparse feature maps. The SENet channel attention module [183] weighed channel importance, determining optimal channels for transmission based on attention weights and the uniqueness of semantic information. A learning-based Irrelevant Feature Remover (IFR) further reduced the transmitted feature map size by replacing extraneous pixels with zeros. This approach significantly reduced data transmission size compared to solutions relying solely on feature encoding.
@@
Another category of learning-based approaches, the selective-region mechanism, focuses on selecting CPM information by considering relevant spatial areas. These methods ensure that only pertinent information within appropriate communication regions is transmitted, avoiding bandwidth waste. For instance, [168] employed a bounding box matching module and strategic critical point selection to choose features enclosed within bounding box proposals for transmission. In contrast, MRSR [170] introduced the Entropy-CS, a trainable region-wise communication selection module, leveraging entropy to distinguish informative regions and select suitable regions for transmission based on resolution levels. GevBEV [171] identifies areas requiring supplemental information for the ego by utilizing evidential BEV maps. This method employs evidential deep learning, incorporating Dirichlet evidence [184], to quantify point-based classification uncertainty within BEV maps. This quantification aids in pinpointing the most relevant regions for communication.

### v4

@@ IV-B1 Cooperative Agent Selection
Selecting the right collaborators is crucial because inappropriate candidates can cause disruptions and disturbances that could harm overall performance.
The attention mechanism, recognized for its efficacy in improving model efficiency by focusing on relevant data, offers promise in selecting suitable agents for the ego CAV. In this context, a multi-stage handshake communication mechanism was proposed in Who2com [115], employing the general attention mechanism [152] to select the appropriate collaborators. This mechanism computes matching scores among agents, selectively engaging the most pertinent agents to optimize bandwidth usage. Based on Who2com, When2com [116] introduced scaled general attention to ascertain optimal communication timings, enabling the ego agent to initiate communication solely under conditions of insufficient information, thereby conserving network resources efficiently. Although this multi-stage communication process proves effective, its implementation and optimization in dynamically varying environments pose significant challenges. To address these challenges, [117] proposed a dynamic communication mechanism that allows vehicles to dynamically determine which infrastructures to communicate with based on the learned attention weights, thus striking a balance between detection accuracy and communication spectrum resources.
@@
To minimize energy consumption, the data sharing scheduling problem can be framed as a variant of the Multi-Armed Bandit (MAB) problem [153], factoring in perception performance, time-varying wireless channels, and power consumption. In this regard, an online learning-based algorithm known as the Adaptive Volatile Upper Confidence Bound (AVUCB) [118] was proposed to schedule the most advantageous vehicle, offering an enhanced view while adhering to communication bandwidth constraints. However, it does not account for the high mobility of cooperative vehicles. To address this issue, the Mobility-Aware Sensor Scheduling (MASS) algorithm [119] formulated decentralized CP scheduling as a Restless-MAB (RMAB) problem [154]. In the RMAB context, rewards continually evolve, requiring each vehicle to continuously learn about its surroundings while selecting actions that empirically maximize rewards. When applied to a practical SUMO trace, the MASS algorithm demonstrated the ability to enhance overall perception gain without incurring additional costs associated with frequent meta-information exchanges.
@@
Cooperative agent selection can also be likened to a Multi-agent Path Finding (MAPF) problem, which plans paths for multiple agents [155]. Inspired by MAPF, a selective communication algorithm [120] was proposed to determine the subset of cooperative vehicles based on information gains from connected vehicles. The ego-vehicle estimates information gains by comparing its detected objects’ positions to received 2D centre positions from other connected vehicles. The vehicles associated with the highest information gain are selected to form the communication scope. In contrast to random selection, this algorithm achieved higher success rates with minimal additional communication costs.
@@
All the aforementioned methods primarily focused on spatial correlations, overlooking the significance of temporal dependencies in collaborator selection. To address this gap, IoSI-CP [121] capitalized on the importance of semantic information from both temporal and spatial dimensions to select optimal collaborators. This approach leveraged a Graph Neural Network (GNN) to capture inherent correlations between the ego vehicle and its neighbouring agents. Agents with weights of zero or lower were considered ineligible for collaboration, which eliminates unsuitable collaborators responsible for noise. Nevertheless, this approach relied on one-dimensional weights that may not fully capture the intricacies addressed by high-dimensional representation methods. InterCoop [122] computes an interaction score for each vehicle by analyzing the geometric properties of the road network in conjunction with the temporal dynamics of the trajectories, prioritizing those most relevant to the prevailing driving conditions.
@@ IV-B2 Perception Information Selection
The selection of targeted information for the ego vehicle is another essential factor in efficient CP. Recently, a growing focus has been on the selection of cooperative message content to improve communication efficiency. A number of methods have emerged for selecting or generating CPMs intended for transmission. These methods can be broadly categorized into rule-based, distance-based, and learning-based approaches, each rooted in distinct mechanisms and theoretical principles.
@@
The rule-based approach to CPM generation/selection entails designing rules that dictate which aspects of detected object information should be included in a CPM. For example, the work in [123] proposed a rule that transmits information about an object’s mode only if the object is turning or accelerating. Similarly, in [124], visual information from detected vehicles was used to determine their communication IDs, which are then used to integrate CPMs with local perception information. All detected vehicles that have exchanged CAM messages will not be considered in future CPMs. Standardization efforts, led by the European Telecommunications Standards Institute (ETSI), have also introduced CPM generation rules, specifying what information CPMs should include and when to generate new messages. The work in [125] and [126] has proposed improved algorithms for these rules to filter out redundant information about detected objects that nearby vehicles have recently transmitted. While rule-based methods are straightforward to implement, they rely primarily on changes in the position and speed of detected objects.
@@
Distance-based methods prioritize data based on distance metrics, considering the spatial location relative to the ego vehicle’s onboard sensors to measure data point significance. For instance, the Augmented Informative Collaborative Perception (AICP) system [127] employed a fitness sorting algorithm based on Mahalanobis distance to select crucial information at the application level, effectively filtering out less relevant packets with minimal delay. In a separate approach, Dynamic Feature Sharing (DFS) [128] identified relevant feature data for sharing through Random-K priorities based on Manhattan distance. DFS incorporated early, intermediate, and late fusion strategies within the CP system to optimize dynamic node engagement.
@@
While distance-based methods emphasize the transmission of feature cells or data points close to the ego vehicle, they may only partially meet the ego vehicles’ perception information requirements in real-world scenarios. In contrast, learning-based approaches leverage the entirety of perception data to intelligently determine information selection for sharing. In BM2CP [129], a modality-guided collaboration was designed to selectively share the most critical multi-modal features across agents, thereby promoting communication efficiency. This was achieved by generating a preference threshold mask to filter BEV features. Deep Reinforcement Learning (DRL) allows agents to learn policies through interactions with their environment, maximizing cumulative rewards [156, 157, 158]. The work [130] introduced a DRL-based Cooperative Perception (DRLCP) scheme, where each CAV intelligently decides what information to transmit based on the context captured by its onboard sensors. DRLCP’s primary objective is to determine whether to transmit or discard data, but it was evaluated on only two road networks without considering the impact of obstructions on vehicular communications.
@@
Another category of learning-based approaches, the selective-region mechanism, focuses on selecting CPM information within appropriate communication regions. For instance, [131] employed a bounding box matching module and strategic critical point selection to choose features enclosed within bounding box proposals for transmission. In contrast, UMC [132] introduced the Entropy-CS, a trainable region-wise communication selection module, leveraging entropy to distinguish informative regions and select suitable regions for transmission based on resolution levels. GevBEV [133] used evidential BEV maps to identify areas where the ego vehicle needs additional information, applying evidential deep learning with Dirichlet evidence [159] to quantify point-based classification uncertainty, thereby highlighting the most relevant regions for communication.
Foreground regions contain more informative content than background regions. Based on this principle, the IFTR [134] employed a message selection and feature map reconstruction module to share the most relevant visual features from the foreground regions. In CodeFilling [135], the information-filling-driven message selection was designed to determine the ego vehicle’s perceptual needs by solving a local optimization problem based on the information score map gathered from other agents. An optimized selection matrix is produced to dictate which spatial regions should be prioritized for transmission. Spatial selection transmits only the informative regions identified by spatial confidence maps, but as bandwidth constraints become more stringent, this approach may lead to the potential loss of objects [136]. In contrast, V2X-PC [136] introduced a Point Cluster Packing module to control bandwidth usage by sampling important points. The point cluster representation preserves object features during message packing, enables efficient message aggregation regardless of collaboration range, and allows for explicit communication of object structure information.
Attention mechanism enables a single agent to intelligently identify the most relevant perception information from multiple agents. For example, the pillar attention encoder (PAE) [137] extracted the attention value of each feature as an importance indicator and selected the most important features for transmission based on communication bandwidth. In ActFormer [138], an active selection network was proposed to assign each query with an interest score, which indicates the relevance and significance of each feature for the ego’s perception. This method actively selects the most relevant sensory data from multiple robots based on spatial knowledge without relying on the sensory measurements themselves, which improves the efficiency of multi-agent object detection.
IV-B3 Agent and Information Selection
Methods that consider both cooperative agent and information selection jointly can offer a more balanced approach to optimizing precision and bandwidth in CP. One such approach, Where2comm [139], introduces a spatial-confidence-aware CP framework that employs spatial confidence maps generated by each agent to identify perceptually critical regions in the feature map. These critical features are compactly packaged into messages and shared via a sparsely connected communication graph. A communication graph construction strategy is employed to selectively determine communication partners, thereby minimizing unnecessary bandwidth usage. Furthermore, only non-zero features and their corresponding indices are transmitted. Experimental results indicate that Where2comm effectively facilitates mutually beneficial cooperation by sharing spatially sparse yet crucial features through V2X communication, optimizing both perception accuracy and communication efficiency.
IV-B4 Perception Information Compression
Compression techniques can effectively prioritize and retain the most critical features of perception data, ensuring that even with reduced data size, the essential information required for perception tasks is preserved. For example, V2VNet [140] utilized a variational image compression algorithm [160] to compress feature maps. In [141], the feature maps from each vehicle are compressed through quantization to reduce communication costs. This on-demand sharing mechanism enabled efficient cooperation among connected vehicles by selectively sharing compressed feature maps but ignored the integration of spatial-temporal information. CPSC [142] performed feature-level compression by focusing on critical regions of perceptual information in the spatial-temporal domain while adapting the data transmission strategy based on network conditions.
Feature encoding using Convolutional Neural Networks (CNNs) is a widely adopted method for feature compression. FS-COD [143] employed a filter at the last layer of the CNN-based feature extractor to control the size of the shared data between cooperative vehicles. DiscoNet [144] used a $1\times 1$ convolutional autoencoder [161] to compress feature maps along the channel dimension, while V2X-ViT [145] applied multiple $1\times 1$ convolutional filters for the same purpose. COOPERNAUT [146] introduced a Point Encoder using the Point Transformer [162] to compress point clouds into compact representations, filtering partially processed data to balance perception accuracy and communication load. Similarly, CenterCoop [147] encoded informative cues from the entire BEV context into compact center representations, significantly reducing communication costs by transmitting these compact center features as queries instead of the entire feature map.
While CNN-generated feature maps offer potential data volume reduction, their dimensions often exceed transmission feasibility within current V2X technology scopes [148]. In [148], the SENet channel attention module [163] was used to weigh channel importance, determining optimal channels for transmission based on attention weights and the uniqueness of semantic information. EMIFF [149] introduced a Feature Compression module with channel and spatial compression blocks for transmission efficiency. Although these approaches effectively reduced the necessary transmission bandwidth, they did not account for the heterogeneity of transmitted information resulting from configuration disparities among agents [164]. To address this issue, What2comm [150] introduced a communication mechanism based on feature decoupling that strategically minimizes bandwidth usage by eliminating background noise to isolate sparse yet essential exclusive features and by compressing common features along the channel dimension.
The aforementioned compression methods have uniformly applied the same compression ratio to all CAV data, but this fairness approach fails to account for the varying contributions of CAVs to perception during data fusion [165]. SmartCooper [151] optimized vehicle connectivity while considering communication constraints, using a learnable encoder to dynamically adjust the compression ratio based on channel state information and employing a judger mechanism to filter out detrimental image data.
IV-C V2X Communication
CP depends on V2X communication technologies to facilitate real-time information exchange among vehicles, infrastructure, and other road users. Dedicated Short-Range Communications (DSRC) and Cellular V2X (C-V2X) are two key technologies that support the low-latency and scalable communication needed in CP systems. Each technology has distinct features and challenges, which are explored below, followed by an analysis of how these limitations impact CP performance and the primary development bottlenecks in V2X communication. A comparison of the key technical differences between DSRC and C-V2X is presented in [166].
IV-C1 DSRC
DSRC was designed specifically for vehicle communication, adapting Wi-Fi technology into the IEEE 802.11p protocol [167]. Operating in the 5.9 GHz band, DSRC enables fast, two-way communication over short distances among vehicles, infrastructure, and other road users. Its decentralized design enables direct data exchange between vehicles without relying on a central network, minimizing latency [167], but the communication range is limited to less than 1km, and the link capacity is less than 10 Mbps [168]. Thus, DSRC can be used to implement Applications that require minimum latency, like collision warnings [166].
IV-C2 C-V2X
C-V2X, standardized by the 3rd Generation Partnership Project (3GPP), enables broad V2X interactions encompassing V2V, V2I, V2N, and V2P connections. It operates in two modes defined in 3GPP Releases 14 through 16: Direct Communication (PC5 interface) and Network Communication (Uu interface) [169, 170, 171]. The PC5 interface enables direct communication between vehicles and nearby infrastructure in the 5.9 GHz band, ideal for low-latency, localized applications such as crash avoidance and real-time safety alerts. The Uu interface, meanwhile, connects vehicles to broader V2N applications through LTE or 5G networks, supporting long-range data exchange for services like traffic management and infotainment, particularly for non-latency-sensitive applications.
IV-C3 Impact of V2X on CP
V2X communication enables data sharing, which is essential for CP, but it faces limitations that restrict CP’s full potential. Key challenges impacting CP performance are bandwidth and spectrum constraints, as well as latency and synchronization issues.
Bandwidth and Spectrum Constraints
CP demands real-time, high-quality data, often involving high-resolution inputs from sensors such as LiDAR and radar, which places significant data rate requirements on communication systems. Achieving high data rates typically requires large bandwidth, stable transmission channels, and precise channel detection and estimation. DSRC operates within the 5.9 GHz frequency band with a total bandwidth of 75 MHz, divided into seven 10 MHz channels ($5.850-5.925$ GHz). However, due to high vehicle mobility, dynamic environmental changes, and multipath delay spread, DSRC’s communication channels are highly time- and frequency-selective [172]. Studies estimate that up to 50% of the coherence bandwidth is in the order of 1 MHz, with coherence times as short as 0.2 ms [172], limiting reliable transmission times and increasing dropout rates. Similar constraints affect C-V2X, which operates on a narrower 30 MHz spectrum ($5.895-5.925$ GHz) [173]. The PC5 mode of C-V2X offers an advantage by utilizing a blind hybrid automatic repeat request (HARQ) technique that combines turbo coding (providing better performance than the convolutional coding used in DSRC) with automatic repeat request error control and soft-combining [174].
Addressing these challenges requires improved network planning and advanced channel detection and estimation techniques. For DSRC, [175] proposes an analytical model to assess key metrics such as Channel Busy Ratio and Packet Delivery Ratio over distance, facilitating optimized network planning. Additionally, [176] enhances IEEE 802.11p through an OFDM-IM system that strategically positions pilot signals and leverages inactive channels to increase detection accuracy, while [177] introduces a low-complexity, high-speed deep neural network-augmented least square algorithm, effectively reducing model size and latency while improving accuracy under varied noise levels and channel conditions. For C-V2X, [178] proposes a balanced resource allocation strategy that evenly distributes resources across subframes, decreasing packet loss and enhancing data throughput. Similarly, [179] suggests a pilot design with zero-correlation zone sequences and a time-domain channel estimation approach to mitigate co-channel interference in C-V2X systems.
Latency and Synchronization
CP requires low latency and precise timing to ensure synchronized data fusion across vehicles. Although DSRC and C-V2X are both designed to minimize latency, high-density environments, network congestion, and physical obstructions can introduce delays. Studies indicate that DSRC and C-V2X can support applications with an end-to-end latency of approximately 100 milliseconds under moderate vehicular density [180]. However, CP demands stricter latency and high-quality service standards to function effectively. Additionally, synchronization challenges arise when fusing data from multiple agents, often leading to temporal mismatches that degrade CP accuracy. Both DSRC and C-V2X utilize the Global Navigation Satellite System (GNSS) for time synchronization, but DSRC operates with looser timing and asynchronous communication, while C-V2X adheres to stricter time standards [166]. Despite these efforts, network congestion can still impact latency and synchronization reliability.
To address congestion in DSRC, [181] proposes adjustments to vehicle settings based on information from supplementary channels, enabling lead vehicles to monitor hidden nodes in a group. Furthermore, [182] presents a congestion-aware forwarding scheme that optimizes multi-hop forwarding reliability and reduces latency under heavy network load. In C-V2X, [179] introduces a rate control mechanism that allows power control to actively contribute to congestion management, while [183] suggests a DQN-based adaptive congestion control algorithm, where a central controller trains and distributes the policy, improving QoS compliance over ETSI’s Distributed Congestion Control in dense traffic scenarios. For synchronization, [184] offers a decentralized clock synchronization protocol to enhance DSRC’s resilience against GPS spoofing and improve reliability when GNSS signals are unavailable. Meanwhile, [185] proposes a synchronization signal design that uses generalized reciprocal sequences for mmWave C-V2X. This approach enhances OFDM-based V2V synchronization by achieving high detection accuracy, minimizing interference, and reducing computational complexity.
IV-D Cooperative Information Alignment
Cooperative information alignment is crucial for CP systems in autonomous driving to ensure that shared data from different sources are accurately represented within a common coordinate system [186]. This process is essential for fusing information from multiple agents to achieve a unified understanding of the environment.
IV-D1 Data Alignment
Each collaborative agent typically operates within its local coordinate system, which may differ in both position and orientation relative to the ego agent. Data shared from cooperative agents, whether raw data, intermediate features, or perception outcomes, is initially referenced in each agent’s local frame. To align the cooperative data from a collaborative agent (Agent B) with the ego vehicle (Agent A), the messages transmitted must include the sender’s position and its orientation. Upon receiving data and pose information from the collaborative agent, the ego agent calculates a relative transformation matrix, $T_{\mathrm{B\rightarrow A}}$, to map the incoming data into the ego agent’s coordinate frame.
To compute $\mathbf{T}_{\mathrm{B\rightarrow A}}$, it is essential first to establish the transformation matrix $\mathbf{T}$ that projects each agent’s local coordinate system onto the global coordinate frame. This matrix incorporates both rotation and translation components, determined by each agent’s global position $[x,y,z]$ and orientation angles $\alpha$, $\beta$, and $\gamma$ (representing yaw, pitch, and roll, respectively). The transformation matrix $\mathbf{T}$ is constructed as follows:
$\mathbf{T}=\begin{bmatrix}r_{11}&r_{12}&r_{13}&x\\ r_{21}&r_{22}&r_{23}&y\\ r_{31}&r_{32}&r_{33}&z\\ 0&0&0&1\end{bmatrix},$ (9)
where $r_{ij}$ ($i,j\in\{1,2,3\}$) denotes the entries of the 3D rotation matrix $\mathbf{R}$, which can be calculated as:
$\boldsymbol{R}=\boldsymbol{R}_{z}(\alpha)\cdot\boldsymbol{R}_{y}(\beta)\cdot\boldsymbol{R}_{x}(\gamma),$ (10)
where $\boldsymbol{R}_{z}(\alpha)$, $\boldsymbol{R}_{y}(\beta)$, and $\boldsymbol{R}_{x}(\gamma)$ are the basis rotation matrices around the $z$, $y$, and $x$ axes, respectively:
$\boldsymbol{R}_{z}(\alpha)=\begin{bmatrix}\cos(\alpha)&-\sin(\alpha)&0\\ \sin(\alpha)&\cos(\alpha)&0\\ 0&0&1\end{bmatrix},$ (11)
$\boldsymbol{R}_{y}(\beta)=\begin{bmatrix}\cos(\beta)&0&\sin(\beta)\\ 0&1&0\\ -\sin(\beta)&0&\cos(\beta)\end{bmatrix},$ (12)
$\boldsymbol{R}_{x}(\gamma)=\begin{bmatrix}1&0&0\\ 0&\cos(\gamma)&-\sin(\gamma)\\ 0&\sin(\gamma)&\cos(\gamma)\end{bmatrix}.$ (13)
Based on (9), the transformation matrices $\mathbf{T}_{\mathrm{A}}$ and $\mathbf{T}_{\mathrm{B}}$ for agents A and B can be derived from their respective poses. To find the relative transformation matrix $\mathbf{T}_{\mathrm{B}\rightarrow\mathrm{A}}$, multiply the inverse of $\mathbf{T}_{\mathrm{B}}$ by $\mathbf{T}_{\mathrm{A}}$:
$\mathbf{T}_{\mathrm{B}\rightarrow\mathrm{A}}=\mathbf{T}_{\mathrm{B}}^{-1}\cdot\mathbf{T}_{\mathrm{A}}.$ (14)
Each data point from agent B’s perspective is represented as $\mathbf{p}_{\mathrm{B}}=[a,b,c]$. Converting these 3D points to homogeneous coordinates by appending a $1$ as the fourth element yields $\mathbf{p}_{\mathrm{B}}=[a,b,c,1]^{\mathrm{T}}$. To align each data point with the ego vehicle’s frame, multiply it by $\mathbf{T}_{\mathrm{B}\rightarrow\mathrm{A}}$:
$\mathbf{p}=\mathbf{T}_{\mathrm{B}\rightarrow\mathrm{A}}\cdot\mathbf{p}_{\mathrm{B}},$ (15)
where $\mathbf{p}$ is the transformed point in the ego agent’s coordinate frame. This alignment enables the direct fusion of incoming data from other agents with the ego’s data, as all points are now within the same coordinate system.
IV-D2 Alignment for Pose Errors
Accurate relative transformation matrices are foundational for ensuring alignment in a shared coordinate system. However, these matrices depend on sensor measurements from posture detection systems such as Global Positioning System (GPS) and Inertial Measurement Unit (IMU), which are prone to noise and accuracy limitations, potentially leading to misalignment and reduced perception performance [187].